If Sally can paint a house in 4 hours, & John can paint the same house in 6 hours, how long will it take for both of them to paint the house together?
Found 2 solutions by nerdybill, J2R2R:
Answer by nerdybill(7384)

If Sally can paint a house in 4 hours, & John can paint the same house in 6 hours, how long will it take for both of them to paint the house together?
.
Let x = time (hours) it takes for both
then
x(1/4 + 1/6) = 1
multiplying both sides by 12:
x(3 + 2) = 12
x(5) = 12
x = 12/5 hours
x = 2.4 hours
or
x = 2 hours and 24 minutes

Answer by J2R2R(94)

Sally does 1/4 of the house in an hour and John does 1/6 of the house in an hour, so both they do 1/4 + 1/6 = 5/12 of the house in an hour.

Which means they take 12/5 hours or 2.4 hours to paint the house.

Check:
Sally’s proportion of the house is 2.4/4 = 3/5
John’s proportion of the house is 2.4/6 = 2/5
Combined they give 2/5 + 3/5 = 1 which is the full house.

You can see Sally does 3/2 times as much as John because she works 3/2 times as fast and therefore takes 2/3 of the time to do the same job (4 hours as opposed to 6 hours).

Hope this makes sense.